The crank and connecting rod of a theoretical steam engine are 0.5 m and 2 m long respectively. The crank makes 180 r.p.m. in the clockwise direction. When it has turned 45° from the inner dead centre position, determine : 1. Velocity and acceleration of the piston 2. Angular velocity and angular acceleration of the connecting rod 3. Velocity and acceleration of point E on the connecting rod 1.5 m from the gudgeon pin 4. Position and linear velocity of any point G on the connecting rod which has the least velocity relative to crank shaft 


1
Expert's answer
2021-06-15T06:07:01-0400

"N _{\nBO} =180 r.p.m. \\implies \u03c9_{ \nBO}\n\u200b\n =2\u03c0\u00d7 \\frac{180}{60}=18.852rad\/s"

"v \n_{BO}\n\u200b\n =v{ \nB}\n\u200b\n =\u03c9 _{\nBO}\n\u200b\n \u00d7OB=18.852\u00d70.5=9.426m\/s"

By measurement, we find that velocity of piston P, vP ​= vector op=8.15m/s

vPB = vector bp=6.8m/s

Since the length of connecting rod PB is 2 m, therefore angular velocity of the connecting rod,

"\u03c9 _{\nPB}\n\u200b\n = \\frac{v_{PB}}{PB} = = \\frac{6.8}{2}\n\u200b\n =3.4rad\/s( Anticlockwise )"

By measurement, we find that velocity of point E, "v _\nE\n\u200b\n = vector oe=8.5m\/s"

"\\frac{BE}{BP}\n\u200b\n = \\frac{\nbp}{\nbe} \\space \n\u200b\n or \\space be= \n\\frac{BE\u00d7bp}{BP}\n\u200b"

We know that velocity of rubbing at the pin of crank-shaft

"= \\frac{d_ \nO}{2}\n\n\u200b\n \n\u200b\n \u00d7\u03c9_{ \nBO}\n\u200b\n =\n2\n0.05\n\u200b\n \u00d718.85=0.47m\/s"

Velocity of rubbing at the pin of crank

"= \n\\frac{d _\nB}{2}\n\n\u200b\n \n\u200b\n (\u03c9 _{\nBO}\n\u200b\n +\u03c9_{\nPB}\n\u200b\n )= \n\\frac{\n0.06}{2}\n\u200b\n (18.85+3.4)=0.6675m\/s"

ωBO ​ is clockwise and ωPB ​ is anticlockwise and velocity of rubbing at the pin of cross-head

"= \n\\frac{\nd_ \nC}{2}\n\u200b\n \n\u200b\n \u00d7\u03c9 _{\nPB}\n\u200b\n = \\frac{\n0.03}{2}\n\u200b\n \u00d73.4=0.051m\/s"

∴At the cross-head, the slider does not rotate and only the connecting rod has angular motion.

By measurement, we find that vector bg = 5 m/s

The position of point G on the connecting rod is obtained as follows

"\\frac{b_g}{b_p}\n\n\u200b\n = \\frac{BG}{BP}\n\u200b\n or BG= \\frac{bg}{bp}\n\n\n\u200b\n \u00d7BP= \n\\frac{5}{6.8}\n\u200b\n \u00d72=1.47m"

By measurement, we find that the linear velocity of point G, vG ​= vector og=8m/s